(1 point) Solve the equation -4x - 7y - 92 -13. X ܕ ܠ ܕ +S +t
Answer #1

Solution: Giren egucetion -47-7y -9Z 213 o claim. To find solution of equoedion from egucehon (0) -456 = 13 +74 +92 x= -13 -